OpenStudy (anonymous):

simplify radical √288

OpenStudy (anonymous):

\[12\sqrt{2}\]

OpenStudy (anonymous):

show work!!

OpenStudy (anonymous):

Because that is simplified down to \[\sqrt{144*2}\] = \[12\sqrt{2}\]

OpenStudy (anonymous):

Sheesh, I was typing.

OpenStudy (anonymous):

so is 12radical 2 the answer??

OpenStudy (anonymous):

nice answer